OpenStudy (anonymous):

write a slope-intercept equation for the line that passes through (-8,-11) and is perpendicular to -2x-7y=-14

OpenStudy (anonymous):

-2x-7y=-14 y=(-2/7)x+2 let s=gradient of y=-2/7 let t=gradient of line that passes through (-8,-11) and is perpendicular to -2x-7y=-14 s*t=-1 (-2/7)*t=-1 t=7/2 Y=tX+C , (-8,-11) -11=(7/2)(-11) + C C=55/2 Y=(7/2)X+ 55/2